### v3

# Seeding neural network quantum states with tensor network states

Abstract
We find an efficient approach to approximately convert matrix product states (MPSs) into restricted Boltzmann machine wave functions consisting of a multinomial hidden unit through a canonical polyadic (CP) decomposition of the MPSs. This method allows us to generate well-behaved initial neural network quantum states for quantum many-body ground-state calculations in polynomial time of the number of variational parameters and systematically shorten the distance between the initial states and the ground states while increasing the rank of the CP decomposition. We demonstrate the efficiency of our method by taking the transverse-field Ising model as an example and discuss possible applications of our method to more general quantum many-body systems in which the ground-state wave functions possess complex nodal structures.
I Introduction
Solving quantum many-body problems is one of the most challenging tasks in modern physics, and tensor network states have been widely used to efficiently represent quantum many-body states in recent years. Matrix product states (MPSs) [1, 2, 3, 4, 5, 6, 7, 8, 9, 10], often specialized in one spatial dimension, and their generalizations to higher dimensions, such as projected entangled pair states (PEPSs) [11, 12, 9, 13, 10, 14] and tree tensor networks [15, 16, 17, 18], have been successfully applied to various quantum many-body problems in low-dimensional quantum systems by keeping the entanglement entropy of the wave function as low as possible. The number of variational parameters in tensor network states remains relatively small and grows only polynomially with the number of sites in most of quantum many-body systems.
Recently, neural network quantum states (NNQSs) have been proposed as a new class of variational wave functions for quantum many-body systems [19, 20, 21, 22, 23, 24, 25, 26]. One of the basic NNQSs is the restricted Boltzmann machine (RBM) wave function [19, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38]. In contrast to tensor network states, NNQSs try to represent quantum many-body states by overparameterizing variational wave functions with neural networks. Because of the flexibility of the network structure and the large number of variational parameters, NNQSs can, in principle, represent arbitrary quantum many-body states even with the volume-law entanglement entropy [39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50], irrespective of the spatial dimensionality of quantum systems [51, 52, 53].
Despite the advantages in representing quantum many-body states with NNQSs, the large number of variational parameters in NNQSs often complicates the optimization of the wave function. Optimized NNQSs may become trapped in local minima within the energy landscape, leading to inaccurate energy estimations. Such difficulties are sometimes mitigated by the use of initial states that are sufficiently close to the quantum states of interest. For best performance, NNQSs ought to be seeded with a well-prepared initial configuration. However, there have been few studies on generating initial states for NNQSs that are well-suited for optimization toward ground states.
Figure 1: Schematic figure of the conversion from MPSs to the RBM wave function through the CP decomposition. (a) MPS representation of tensor $T$. (b) CP decomposition of tensor $T$, which is efficiently computed from the MPS representation. (c) RBM wave function with a multinomial hidden unit, which is equivalent to the CP decomposed tensor.
In this paper, we find that suitable initial states of NNQSs can efficiently be prepared by tensor network states. Previously, there have been numerous studies on constructing tensor network states, MPSs in particular, from RBM wave functions [47, 54, 55, 56, 48, 49, 50, 57, 58, 59, 60, 61, 62, 63, 64]. On the contrary, although it is widely believed that NNQSs have the same or higher expressive power than tensor network states [43, 25], examples of generating RBM wave functions from MPSs have been limited to special cases [47, 54, 55, 56, 48, 57, 58, 49, 59, 60, 61]. Such examples include RBM wave functions from primitive tensor network states, essentially described by the so-called stabilizer state [65] such as the Greenberger-Horne-Zeilinger (GHZ) state (e.g., $|\mathrm{GHZ}\rangle\propto|000\rangle+|111\rangle$ for three qubits) [66, 67] or more general (in a sense that it is a non-stabilizer state) but still a simple one such as the W state (e.g., $|\mathrm{W}\rangle\propto|001\rangle+|010\rangle+|100\rangle$ for three qubits) [68]; otherwise, resulting Boltzmann machine wave functions require more than one hidden layer [50, 61, 49, 59, 60, 62, 63, 64], thereby losing the structural simplicity inherent to shallow network configurations.
We explore the possibility of generating RBM wave functions from MPSs in a more general setting. To this end, instead of directly transforming MPSs into RBM wave functions, we introduce intermediate tensor network states, which can efficiently be obtained by the canonical polyadic (CP) decomposition [69, 70, 71, 72, 73, 74, 75, 76, 77] of arbitrary MPSs for a given rank of the decomposition (see Fig. 1). The CP decomposed tensor network state is found to be equivalent to the RBM wave function having a single multinomial hidden unit. Therefore, when we are able to prepare accurate MPSs of a given quantum system using conventional tensor network methods, we can easily generate equivalent RBM wave functions as initial states of NNQSs.
This paper is organized as follows: In Sec. II, we introduce our method for approximately converting MPSs into RBM wave functions consisting of multinomial hidden units by the CP decomposition of the MPSs. The computational complexity of our method scales polynomially with the number of variational parameters when the practical rank of the CP decomposition is assumed to be known and is fixed by the user. In Sec. III, we demonstrate the efficiency of our method by taking the transverse-field Ising model as an example. We first examine systems with open boundary conditions and further test systems with periodic boundary conditions using the initial states generated by our method under open boundary conditions. Finally, in Sec. IV, we summarize our results and discuss possible applications of our method to more general quantum many-body systems in which the ground states possess complex nodal structures.
II Methods
In this section, we introduce our method for approximately converting MPSs into RBM wave functions consisting of multinomial hidden units by the CP decomposition of the MPSs.
II.1 MPS representation of tensors
When the order of a tensor $T$ increases, the number of elements in $T$ grows exponentially large; eventually, it is not possible to store the full tensor $T$ in memory. Such a tensor having a large number of axes often appears in calculating the ground state of quantum many-body systems with a large number of sites $n$ using the exact diagonalization method. The wave function is represented as
$|\Psi\rangle=\sum_{\{s_{i}\}}T_{s_{1},s_{2},\dots,s_{n}}|s_{1},s_{2},\dots,s_{n}\rangle,$ (1)
where $s_{i}$ denotes the physical index at site $i$. In such cases, we can use the MPS representation [1, 2, 9, 3, 4, 5, 10, 6, 7, 8] of the tensor, which can be more efficiently obtained by the density matrix renormalization group (DMRG) method [2] than the exact diagonalization method.
The MPS representation of tensor $T$ only requires parameters of the form $G^{(i)}\in\mathbb{C}^{D_{i-1}\times d_{i}\times D_{i}}$, where $d_{i}$ denotes the physical bond dimension at site $i$ and $D_{i}$ ($D_{0}=D_{n}=1$) are the virtual bond dimensions. For a spin-$1/2$ system, a physical spin index $s_{i}$ takes the value $\uparrow$ or $\downarrow$, corresponding to $d_{i}=2$. This structure allows for an efficient approximation of high-dimensional tensors, particularly when the entanglement entropy in the system is low. By contracting these matrices $G^{(i)}$ for each $s_{i}$} sequentially, the full tensor $T_{s_{1},s_{2},\dots,s_{n}}\in\mathbb{C}^{d_{1}\times d_{2}\times\cdots\times d_{n}}$ can be reconstructed as
$T_{s_{1},s_{2},\dots,s_{n}}=\sum_{\{\alpha_{i}\}}G^{(1)}_{\alpha_{0},s_{1},\alpha_{1}}G^{(2)}_{\alpha_{1},s_{2},\alpha_{2}}\cdots G^{(n)}_{\alpha_{n-1},s_{n},\alpha_{n}}$, (2)
where the internal indices $\alpha_{i}(=1,2,\dots,D_{i})$ are summed over according to the bond dimensions.
II.2 CP decomposition of tensors
We first review the CP decomposition of an order-$n$ tensor $X_{i_{1},i_{2},\dots,i_{n}}\in\mathbb{C}^{d_{1}\times d_{2}\times\cdots\times d_{n}}$, where $d_{i}$ is the dimension of the $i$th axis [69, 70, 71, 72, 73, 74, 75, 76, 77]. The tensor elements are expressed by
$X_{i_{1},i_{2},\dots,i_{n}}=\sum_{r=1}^{R}A^{(1)}_{i_{1},r}A^{(2)}_{i_{2},r}\cdots A^{(n)}_{i_{n},r}$, (3)
where $R$ is the rank of the CP decomposition and $A^{(j)}_{i_{j},r}\in\mathbb{C}^{d_{i}\times R}$ is the $r$th component of the vector corresponding to the $j$th axis. The CP decomposition becomes more accurate as the rank $R$ increases. We discuss the case where the tensor is represented as MPSs in the next section.
One of the most famous algorithms to compute the CP decomposition is the alternating least squares (ALS) method [78, 77, 79]. For simplicity, let us consider the case with $n=3$. Our goal is to find matrices $A$, $B$, and $C$ that minimize the following loss function:
$L(A,B,C)$ $=\sum_{i,j,k}\left|X_{i,j,k}-\sum_{r=1}^{R}A_{i,r}B_{j,r}C_{k,r}\right|^{2}.$ (4)
The ALS method iteratively updates matrices $A$, $B$, and $C$. The initial values of $A$, $B$, and $C$ are often randomly chosen. In the first step, we fix $B$ and $C$ and minimize the loss function with respect to $A$. By taking the derivative of the loss function with respect to $A_{i,r}$ and setting the derivative to zero, we see that the optimal element $A_{i,r}$ is given by
$A_{i,r}=\sum_{j,k}X_{i,j,k}\sum_{r^{\prime}}B^{*}_{j,r^{\prime}}C^{*}_{k,r^{\prime}}\left(\sum_{j^{\prime},k^{\prime}}B^{*}_{j^{\prime},r^{\prime}}B_{j^{\prime},r}C^{*}_{k^{\prime},r^{\prime}}C_{k^{\prime},r}\right)^{-1}.$ (5)
The corresponding matrix form is
$A=X_{(1)}\left(C^{*}\odot B^{*}\right)\left(C^{\dagger}C\circledast B^{\dagger}B\right)^{-1},$ (6)
[…]
$B_{j,r}$ $=\sum_{i,k}X_{i,j,k}\sum_{r^{\prime}}A^{*}_{i,r^{\prime}}C^{*}_{k,r^{\prime}}\left(\sum_{i^{\prime},k^{\prime}}A^{*}_{i^{\prime},r^{\prime}}A_{i^{\prime},r}C^{*}_{k^{\prime},r^{\prime}}C_{k^{\prime},r}\right)^{-1}$ (7)
and the matrix form as
$B$ $=X_{(2)}\left(C^{*}\odot A^{*}\right)\left(C^{\dagger}C\circledast A^{\dagger}A\right)^{-1}.$ (8)
Finally, we fix $A$ and $B$ and minimize the loss function $L(A,B,C)$. The optimal element $C_{k,r}$ is given by
$C_{k,r}$ $=\sum_{i,j}X_{i,j,k}\sum_{r^{\prime}}A^{*}_{i,r^{\prime}}B^{*}_{j,r^{\prime}}\left(\sum_{i^{\prime},j^{\prime}}A^{*}_{i^{\prime},r^{\prime}}A_{i^{\prime},r}B^{*}_{j^{\prime},r^{\prime}}B_{j^{\prime},r}\right)^{-1},$ (9)
and the matrix form is given by
$C$ $=X_{(3)}\left(B^{*}\odot A^{*}\right)\left(B^{\dagger}B\circledast A^{\dagger}A\right)^{-1}.$ (10)
We repeat these three steps until the loss function becomes sufficiently small.
The ALS method can be generalized to the case with $n>3$. The loss function is given by
$L(\{A^{(i)}\})$ $=\sum_{i_{1},i_{2},\dots,i_{n}}\left|X_{i_{1},i_{2},\dots,i_{n}}-\sum_{r=1}^{R}A^{(1)}_{i_{1},r}A^{(2)}_{i_{2},r}\cdots A^{(n)}_{i_{n},r}\right|^{2}.$ (11)
The ALS method iteratively updates matrices $A^{(1)}$, $A^{(2)}$, $\dots$, $A^{(n)}$ by fixing all but one of the matrices and minimizing the loss function with respect to the remaining matrix. The update rule for matrix $A^{(j)}$ is given by
$A^{(j)}$ $=X_{(j)}\left[A^{(n)}\odot\cdots\odot A^{(j+1)}\odot A^{(j-1)}\odot\cdots\odot A^{(1)}\right]$
$~\phantom{=}\cdot\bigl[{A^{(n)}}^{\dagger}A^{(n)}\circledast\cdots\circledast{A^{(j+1)}}^{\dagger}A^{(j+1)}$
$~\phantom{=}\circledast{A^{(j-1)}}^{\dagger}A^{(j-1)}\circledast\cdots\circledast{A^{(1)}}^{\dagger}A^{(1)}\bigr]^{-1}.$ (12)
Here, the centered dot $\cdot$ is the conventional matrix product.
In the ALS method, the rank $R$ is a hyperparameter that determines the accuracy of the CP decomposition. The larger the rank $R$ is, the more accurate the CP decomposition becomes. For an order-$3$ tensor, the sufficient number of the rank in the CP decomposition to completely reproduce the original tensor is $\min(d_{1}d_{2},d_{1}d_{3},d_{2}d_{3})$ [77], which is smaller than the number of elements $d_{1}d_{2}d_{3}$ of the tensor. However, there is no algorithm to determine the rank of a given tensor in general. This problem is known to be NP-hard [80]. Furthermore, the ALS method is not guaranteed to converge to the global minimum because of the nonconvexity of the loss function [81]. In realistic applications, we can try several initial values of factors $A$, $B$, and $C$ and numerically determine the practical rank of the tensor by fitting various CP decomposition results with different trials of rank $R$.
When the order $n$ of the tensor is large, the most time-consuming part of the ALS method is the computation of the products between the matricized tensor and the matrix obtained by the Khatri-Rao product. The dimension of the matricized tensor is $d_{i}\times\prod_{j=1,j\not=i}^{n}d_{j}\geq d_{i}\times(\min_{j\not=i}d_{j})^{n-1}$, which results in the exponential cost of evaluating the products. As we will see in the next section, the MPS representation is more efficient than the original tensor representation in terms of memory and computational cost of the ALS method.
II.3 CP decomposition of MPSs
Even when the number of elements in a tensor increases, the memory for the MPS representation remains manageable. Despite the efficiency of the MPS representation, to the best of our knowledge, there are very few studies on the CP decomposition of MPSs. Several algorithms have been proposed to compute the CP decomposition and its variant, such as the Tucker decomposition, of MPSs very recently [82, 83, 84, 85]; however, the rank of the CP decomposition is often smaller than the bond dimension of the MPS. This limitation makes it difficult to obtain accurate CP decompositions of MPSs, especially for MPSs that are highly entangled and require large bond dimensions to represent the wave function accurately.
Figure 2: (a) CP decomposition of a tensor. A black dot represent the Kronecker delta tensor $\delta$. Each factor $A^{(i)}$ is a $d_{i}\times R$ matrix with $R$ being the rank of the CP decomposition. (b) MPS representation of the CP decomposed tensor. The black dot in (a) can be rewritten as a product of Kronecker delta tensors with smaller orders. The product of the Kronecker delta tensor and factor $A^{(i)}$ can be rewritten as MPS $F^{(i)}$.
Figure 3: Tensors needed for the ALS method for the CP decomposition of MPSs. We show an example of a tensor with the order $n=5$ when updating the $j=3$ element. (a) MPS representation of tensor $X$. (b) Matrix $Y^{(j)}$ obtained by the product between the matricized tensor, written in MPS form $G^{(i)}$, and the matrix obtained by the Khatri-Rao product, written by the Kronecker delta tensors $\delta$ and matrix $A^{(i\not=j)}$. Matrix $Y^{(j)}$ is used to update matrix $A^{(j)}$ during the ALS method. (c) Left matrix $L^{(j)}$ at the $j$th step of the ALS method. (d) Right matrix $R^{(j)}$ at the $j$th step of the ALS method. Matrices $L^{(j)}$ and $R^{(j)}$ are used to construct matrix $Y^{(j)}$.
We propose a straightforward method for computing the CP decomposition of MPSs by using the ALS method. Hereafter, we assume that the rank $R$ of the CP decomposition is fixed and given as a hyperparameter. The input is the MPS representation $G^{(i)}$ of tensor $T$, rather than the full tensor $T$ itself, and the output is matrices $A^{(i)}$ that represent the CP decomposition of the MPSs. Here, the MPS representation $G^{(i)}$ is a $D_{i-1}\times d_{i}\times D_{i}$ tensor for $i=1,2,\dots,n$, whereas the factor $A^{(i)}$ is a $d_{i}\times R$ matrix for $i=1,2,\dots,n$ ($d_{i}=2$ for a spin-$1/2$ system). This procedure can be effectively accomplished because the CP decomposed tensor is already expressed in the form of MPSs. To be more precise, utilizing the Kronecker delta tensor
$\delta_{r_{1},r_{2},\dots,r_{n}}$ $=\begin{cases}1&(r_{1}=r_{2}=\cdots=r_{n})\\ 0&(\text{otherwise}),\end{cases}$ (13)
with $r_{i}=1,2,\dots,R$ for all $i$, we can rewrite the definition of the CP decomposed tensor as
$X_{i_{1},i_{2},\dots,i_{n}}=\sum_{r_{1}=1}^{R}\sum_{r_{2}=1}^{R}\cdots\sum_{r_{n}=1}^{R}\delta_{r_{1},r_{2},\dots,r_{n}}A^{(1)}_{i_{1},r_{1}}A^{(2)}_{i_{2},r_{2}}\cdots A^{(n)}_{i_{n},r_{n}}$. (14)
Because the Kronecker delta tensor $\delta_{r_{1},r_{2},\dots,r_{n}}$ can be expressed as a product of order-$3$ Kronecker delta tensors [86, 87, 61], e.g., $\delta_{r_{1},r_{2},r_{3},r_{4},r_{5}}=\delta_{r_{1},r_{2},r_{3}}\delta_{r_{3},r_{4},r_{5}}$, we can rewrite the above equation as
$X_{i_{1},i_{2},\dots,i_{n}}$ $=\sum_{\{\alpha_{i}\}}F^{(1)}_{\alpha_{0},i_{1},\alpha_{1}}F^{(2)}_{\alpha_{1},i_{2},\alpha_{2}}\cdots F^{(n)}_{\alpha_{n-1},i_{n},\alpha_{n}}$, (15)
$F^{(i)}_{\alpha_{i-1},s_{i},\alpha_{i}}$ $=\delta_{\alpha_{i-1},r,\alpha_{i}}A^{(i)}_{s_{i},r}\quad(i=1,2,\dots,n),$ (16)
where $F^{(i)}$ is the corresponding MPS representation, as shown in Fig. 2. Thus, tensor multiplications in the ALS method can efficiently be performed by the contraction of MPSs.
Let us describe the ALS method specifically for the case of MPSs. We assume that the original tensor $X$ is represented as MPSs $G^{(i)}$, although tensor $X$ is not explicitly given because it is too large to store in memory [see Fig. 3(a)]. We need to evaluate the counterparts of the $d_{i}\times R$ matrix $X_{(1)}(C^{*}\odot B^{*})$ and the $R\times R$ matrix $(C^{\dagger}C\circledast B^{\dagger}B)^{-1}$ in Eq. (6) by contracting the MPSs $G^{(i)}$ and the MPS representations $F^{(i)}$ of factors in the CP decomposition.
The counterpart of the $d_{i}\times R$ matrix $X_{(1)}(C^{*}\odot B^{*})$, as shown in Fig. 3(b), can be evaluated in the following manner: For factors $A^{(j)}$ ($2\leq j\leq n-1$) having the MPS representations $F^{(i)}$, we construct the left and right matrices $L^{(j)}$ and $R^{(j)}$, defined as
$L^{(j)}_{\alpha_{j-1},\tilde{\alpha}_{j-1}}=\sum_{\{\alpha_{i}\},\{s_{i}\},\{\tilde{\alpha_{i}}\}}G^{(1)}_{\alpha_{0}=1,s_{1},\alpha_{1}}F^{(1)*}_{\tilde{\alpha}_{0}=1,s_{1},\tilde{\alpha}_{1}}$
$~\cdot G^{(2)}_{\alpha_{1},s_{2},\alpha_{2}}F^{(2)*}_{\tilde{\alpha}_{1},s_{2},\tilde{\alpha}_{2}}\cdots G^{(j-1)}_{\alpha_{j-2},s_{j-1},\alpha_{j-1}}F^{(j-1)*}_{\tilde{\alpha}_{j-2},s_{j-1},\tilde{\alpha}_{j-1}}$, (17)
$R^{(j)}_{\alpha_{j},\tilde{\alpha}_{j}}=\sum_{\{\alpha_{i}\},\{s_{i}\},\{\tilde{\alpha_{i}}\}}G^{(n)}_{\alpha_{n-1},s_{n},\alpha_{n}=1}F^{(n)*}_{\tilde{\alpha}_{n-1},s_{n},\tilde{\alpha}_{n}=1}$
$~\cdot G^{(n-1)}_{\alpha_{n-2},s_{n-1},\alpha_{n-1}}F^{(n-1)*}_{\tilde{\alpha}_{n-2},s_{n-1},\tilde{\alpha}_{n-1}}\cdots G^{(j+1)}_{\alpha_{j},s_{j+1},\alpha_{j+1}}F^{(j+1)*}_{\tilde{\alpha}_{j},s_{j+1},\tilde{\alpha}_{j+1}}$, (18)
respectively, as shown in Figs. 3(c) and 3(d). Then, we can evaluate the $d_{i}\times R$ matrix $Y^{(j)}$ as
$Y^{(j)}_{s_{j},r}$ $=\sum_{\alpha_{j-1},\tilde{\alpha}_{j-1},\alpha_{j},\tilde{\alpha}_{j}}\delta_{\tilde{\alpha}_{j-1},r,\tilde{\alpha}_{j-1}}L^{(j)}_{\alpha_{j-1},\tilde{\alpha}_{j-1}}G^{(j)}_{\alpha_{j-1},s_{j},\alpha_{j}}R^{(j)}_{\alpha_{j},\tilde{\alpha}_{j}}$. (19)
For factor $A^{(1)}$ $A^{(n)}$], we only need the right (left) matrix to construct matrix $Y^{(j)}$.
The counterpart of the $R\times R$ matrix $(C^{\dagger}C\circledast B^{\dagger}B)$ can be evaluated just as in the original ALS method. We have to calculate the element-wise product of the $R\times R$ matrices n2n-2 times. The corresponding matrix $Z^{(j)}$ is given by
$Z^{(j)}_{r,r^{\prime}}=\prod_{k=1,k\neq j}^{n}\sum_{i_{k}}A^{(k)*}_{i_{k},r}A^{(k)}_{i_{k},r^{\prime}}$. (20)
Finally, we can iteratively update matrix $A^{(j)}$ as
$A^{(j)}_{s_{j},r}$ $=\sum_{r^{\prime}}Y_{s_{j},r^{\prime}}\left[Z^{(j)}\right]^{-1}_{r^{\prime},r}$ (21)
for all $j=1,2,\dots,n$. To safely compute the inverse of matrix $Z^{(j)}$, we add a small constant to the diagonal elements of the matrix $Z^{(j)}$ in practical calculations. We initialize matrices $A^{(j)}$ with random values taken from the standard normal distribution and normalize matrix $A^{(j)}$ in each iteration to avoid numerical instabilities. When the fidelity ($|\langle\mathrm{Orig}|\mathrm{CP}\rangle|^{2}\langle\mathrm{Orig}|\mathrm{Orig}\rangle^{-1}\langle\mathrm{CP}|\mathrm{CP}\rangle^{-1}$) between the MPS representation of the original tensor ($|\mathrm{Orig}\rangle$) and the MPS representation of the CP decomposed tensor ($|\mathrm{CP}\rangle$) is sufficiently high after several iterations, the ALS method gives a good approximation of the CP decomposition.
In general, the most time-consuming part of the ALS method is the computation of the products between the matricized tensor and the matrix obtained by the Khatri-Rao product. When using the MPS representation, this time-consuming part is reduced to the computation of the products among MPSs. The computational cost grows only linearly in the tensor order $n$. This is in contrast to the original ALS method using the full tensor, which requires exponential computational cost in the tensor order $n$.
II.4 RBM wave functions with multinomial hidden units derived from the CP decomposed MPSs
We will see that the CP decomposed MPSs are equivalent to the RBM wave function. Before that, we briefly review the RBM wave function that is used in quantum many-body problems.
In general, the RBM wave function for a spin-$1/2$ system [19],
$|\Psi\rangle=\sum_{\{s_{i}\}}\Psi(\{s_{i}\})|s_{1},s_{2},\dots,s_{n}\rangle,$ (22)
is defined with its wave function amplitude,
$\Psi(\{v_{i}\})$ $=\sum_{\{h_{j}\}}\exp[-E(\{v_{i}\},\{h_{j}\})],$ (23)
$E(\{v_{i}\},\{h_{j}\})$ $=-\sum_{i=1}^{n_{v}}a_{i}v_{i}-\sum_{j=1}^{n_{h}}b_{j}h_{j}-\sum_{i=1}^{n_{v}}\sum_{j=1}^{n_{h}}W_{i,j}v_{i}h_{j}.$ (24)
Here, $a_{i}$, $b_{j}$, and $W_{i,j}$ are the parameters of the RBM. The symbols $n_{v}$ and $n_{h}$ represent the number of the visible and hidden units, respectively. The variable $v_{i}(=s_{i})$ is the index of the $i$th spin and takes the value +1+1 or 1-1, depending on the state of the spin $\uparrow$ or $\downarrow$. For binomial hidden units, the variable $h_{j}$ also takes the value +1+1 or 1-1. Note that, for simplicity, we ignore the normalization factor of the wave function amplitude, which is given by $\sqrt{\langle\Psi|\Psi\rangle}=\sqrt{\sum_{\{s_{i}\}}|\Psi(\{s_{i}\})|^{2}}$.
For the transformation from MPSs to RBM wave functions, we specifically consider the multinomial hidden units that take the states from $1$ to $n_{r}$ [88, 89, 90, 48, 91]. Using the one-hot encoding of the hidden units, we can express the hidden unit $h_{j}$ as an $n_{r}$ component vector,
$h_{j,k}=(0,0,\dots,0,1,0,\dots,0),$ (25)
where only the $k$th component is $1$ and all other components are 0. Then, the virtual energy function is given by
$E(\{v_{i}\},\{h_{j}\})$ $=-\sum_{i=1}^{n_{v}}a_{i}v_{i}-\sum_{j=1}^{n_{h}}\sum_{k=1}^{n_{r}}b_{j,k}h_{j,k}$
$~\phantom{=}-\sum_{i=1}^{n_{v}}\sum_{j=1}^{n_{h}}\sum_{k=1}^{n_{r}}W_{i,j,k}v_{i}h_{j,k},$ (26)
where the sum over $k$ is taken over the states of the hidden unit $h_{j}$. We also add the index $k$ to the parameters $b_{j,k}$ and $W_{i,j,k}$. Defining the parameter
$\theta_{j,k}=b_{j,k}+\sum_{i=1}^{n_{v}}W_{i,j,k}v_{i},$ (27)
we obtain the amplitude in the RBM wave function as
$\Psi(\{v_{i}\})$ $=\sum_{\{h_{j}\}}\exp\left(\sum_{i=1}^{n_{v}}a_{i}v_{i}+\sum_{j=1}^{n_{h}}\sum_{k=1}^{n_{r}}h_{j,k}\theta_{j,k}\right)$ (28)
$=\exp\left(\sum_{i=1}^{n_{v}}a_{i}v_{i}\right)\prod_{j=1}^{n_{h}}\left[\sum_{k=1}^{n_{r}}\exp(\theta_{j,k})\right].$ (29)
We wish to convert the CP decomposed MPSs for $n=n_{v}$ sites, given by
$\Psi(\{s_{i}\})$ $=T_{s_{1},s_{2},\dots,s_{n}}=\sum_{r=1}^{R}A^{(1)}_{i_{1},r}A^{(2)}_{i_{2},r}\cdots A^{(n)}_{i_{n},r}$, (30)
$i_{j}$ $=\frac{1-s_{j}}{2}\quad(j=1,2,\dots,n),$ (31)
into the multinomial RBM wave function. One of such choices is to set the parameters in the following manner:
$a_{i}$ $=0,$ (32)
$n_{h}$ $=1,$ (33)
$n_{r}$ $=R,$ (34)
$\exp(\theta_{j=1,k})$ $=A^{(1)}_{i_{1},k}A^{(2)}_{i_{2},k}\cdots A^{(n_{v})}_{i_{n_{v}},k}$. (35)
By taking the logarithm in the last equation, we obtain
$\theta_{j=1,k}=\sum_{l=1}^{n_{v}}\left(\frac{b_{j=1,k}}{n_{v}}+W_{l,j=1,k}v_{l}\right)=\sum_{l=1}^{n_{v}}\mathrm{Log\,}A^{(l)}_{i_{l},k}$, (36)
where $\mathrm{Log\,}z=\ln|z|+i\arg z$ is the complex logarithm. When $z=0$, we add a small constant to $z$ to calculate $\mathrm{Log\,}z$. Since $v_{l}=\pm 1$ with $i_{l}=0,1$ for a spin-$1/2$ system, formally, it is sufficient to set the parameters as
$\sum_{l=1}^{n_{v}}\frac{b_{j=1,k}}{n_{v}}$ $=\sum_{l=1}^{n_{v}}\frac{\mathrm{Log\,}A^{(l)}_{i_{l}=0,k}+\mathrm{Log\,}A^{(l)}_{i_{l}=1,k}}{2},$ (37)
$W_{l,j=1,k}$ $=\frac{\mathrm{Log\,}A^{(l)}_{i_{l}=0,k}-\mathrm{Log\,}A^{(l)}_{i_{l}=1,k}}{2}$ (38)
for each $l=1,2,\dots,n_{v}$ and $k=1,2,\dots,R$.
By introducing a new site-dependent parameter $b_{l,k}$, corresponding to the parameter $b_{j=1,k}/n_{v}$ in Eq. (37), we finally obtain the RBM wave function for $n$ sites with the rank $R$ as
$\Psi(\{v_{l}\})$ $=\sum_{k=1}^{R}\exp\left(\sum_{l=1}^{n}\theta_{l,k}\right),$ (39)
$\theta_{l,k}$ $=b_{l,k}+W_{l,k}v_{l},$ (40)
with the parameters in the CP decomposed MPSs,
$b_{l,k}$ $=\frac{\mathrm{Log\,}A^{(l)}_{i_{l}=0,k}+\mathrm{Log\,}A^{(l)}_{i_{l}=1,k}}{2},$ (41)
$W_{l,k}$ $=\frac{\mathrm{Log\,}A^{(l)}_{i_{l}=0,k}-\mathrm{Log\,}A^{(l)}_{i_{l}=1,k}}{2}$ (42)
for each $l=1,2,\dots,n$ and $k=1,2,\dots,R$. Therefore, once we obtain the CP decomposition of MPSs, we can convert the MPSs into RBM wave functions immediately. Note that the relation between the CP decomposed tensors and deep Boltzmann machine wave functions with more than one hidden layer is already discussed before [61, 49]. Here, we stick to the simplest RBM wave function without introducing deeper hidden layers. The expressibility of the RBM wave function is extended by allowing the hidden unit to have multinomial values [88, 89, 90, 48, 91].
In general, as the number of hidden units increases, the RBM wave function well approximates the ground state of the quantum many-body system, which is known as the universal approximation theorem [92]. By contrast, as for the present RBM wave function with a single multinomial hidden unit, this theorem does not necessarily hold. However, by definition, the CP decomposition exactly reproduces the original tensor when $R$ is larger than or equal to the CP rank of the tensor [77]; even for $R$ that is smaller than the CP rank, the CP decomposition would nearly reconstruct the original tensor when $R$ is sufficiently large [78, 79]. This fact implies that the RBM wave function with a single hidden unit has a potential to represent the ground state of the quantum many-body system when $R$ is sufficiently large, although $R$ could be exponentially large in the number of sites $n$ in the worst case [93].
II.5 Optimization of RBM wave functions
In the variational Monte Carlo (VMC) simulation, we calculate physical quantities using the Markov chain Monte Carlo sampling over the probability distribution $p(x)=|\langle x|\Psi\rangle|^{2}/\langle\Psi|\Psi\rangle$ with $|x\rangle$ being a real-space spin configuration [94, 95]. We optimize the parameters in the RBM wave function by the gradient-based optimization method. In particular, we use the stochastic reconfiguration (SR) method [96, 97], which is known to be equivalent [27] to the natural gradient method [98, 99]. The derivative of the RBM wave function to calculate the gradient is given by
$\tilde{\theta}_{k}$ $=\sum_{l=1}^{n}\theta_{l,k},$ (43)
$\frac{\partial\ln\Psi(\{v_{i}\})}{\partial b_{l,k}}$ $=\frac{\exp\tilde{\theta}_{k}}{\sum_{k^{\prime}=1}^{R}\exp\tilde{\theta}_{k^{\prime}}},$ (44)
$\frac{\partial\ln\Psi(\{v_{i}\})}{\partial W_{l,k}}$ $=v_{l}\frac{\exp\tilde{\theta}_{k}}{\sum_{k^{\prime}=1}^{R}\exp\tilde{\theta}_{k^{\prime}}}.$ (45)
The right-hand side of the above equations is expressed by the softmax function. To avoid numerical instabilities such as overflow and underflow, we typically calculate a maximum of a real part of $\tilde{\theta}_{k}$ $m=\max_{k}\mathrm{Re\,}\tilde{\theta}_{k}$] and estimate the exponentials in the numerator and denominator by $\exp(\tilde{\theta}_{k}-m)$. The most time-consuming part of the SR method is solving the linear equations with a large number of variational parameters. In general, the computational cost scales quadratically with the number of variational parameters [19].
III Applications to quantum many-body problems: transverse-field Ising model
We examine the applicability of our method by taking the transverse-field Ising model as an example. The Hamiltonian of the one-dimensional transverse-field Ising model is given by
$H=-J\sum_{i}\sigma^{z}_{i}\sigma^{z}_{i+1}-h\sum_{i}\sigma^{x}_{i},$ (46)
where $J$ and $h$ are the strength of the spin-exchange interaction and the transverse field, respectively. For open boundary conditions, the summation over $i$ for the interaction term is taken over $i=1,\ldots,n-1$, where $n$ is the number of sites in the system. For periodic boundary conditions, the summation over $i$ for the interaction term is taken over $i=1,\ldots,n$, and the last site is connected to the first site by the interaction term as $\sigma^{z}_{n+1}:=\sigma^{z}_{1}$.
The ground state of the one-dimensional transverse-field Ising model can be solved analytically using the Jordan-Wigner transformation [100], which maps spin operators onto fermionic creation and annihilation operators. The ground-state energy is given by
$E=-J\sum_{k}\sqrt{1+\lambda^{2}-2\lambda\cos k}\quad(\lambda=h/J).$ (47)
Hereafter, the energy is expressed in the units of $J$. For periodic boundary conditions with even total fermion number, the summation over $k$ is taken over the set of momenta, satisfying
$k=\frac{\pi(2j+1)}{n}\quad(j=0,1,2,\dots,n-1).$ (48)
For open boundary conditions [101, 102], the corresponding $n$ momenta fulfill
$\lambda\sin[(n+1)k]=\sin nk.$ (49)
Hereafter, we first consider the model under open boundary conditions and obtain the ground-state wave function in the MPS representation using the DMRG method. We apply the TeNPy library for the DMRG method [6, 7] and choose the bond dimension that is large enough to represent the true ground-state energy within a desired accuracy, which is smaller than $10^{-9}$ in the units of $J$ in this study. We turn off the parity conservation during the DMRG simulation to purposely keep the wave function primitive. We then apply the CP decomposition to the MPS representation of the ground state. We find that the CP decomposed MPSs are sufficiently close to the ground-state wave function when the rank of the CP decomposition is large. The infidelity between the ground-state MPS and the approximated CP decomposed tensor decreases at most polynomially as the rank of the CP decomposition increases. We then approximately convert the MPS representation into the RBM wave function consisting of a single hidden unit and apply the VMC method to optimize the parameters in the RBM wave function. We demonstrate that the initial state prepared under open boundary conditions can be efficiently used for the simulation under periodic boundary conditions.
Note that the ground state of the present transverse-field Ising model does not have a complex nodal structure. The amplitude of the ground-state wave function can be taken as non-negative according to the Perron-Frobenius theorem [29, 103]. This fact suggests that one can use positive factors for the CP decomposition in this study. However, to analyze the performance of the CP decomposition more generally, we purposely apply the CP decomposition allowing both positive and negative factors. As we see later, the CP decomposition with positive and negative factors also well approximates the ground-state wave function and provides a good initial state having complex amplitudes for the VMC simulation.
III.1 Accuracy of the CP decomposition of MPSs
We calculate two types of errors to evaluate the accuracy of the CP decomposition of MPSs. One is the infidelity $\bar{F}$ between the ground-state MPS ($|\mathrm{MPS}\rangle$) and the approximated CP decomposed tensor ($|\mathrm{CP}\rangle$), defined by
$\bar{F}$ $=1-F,$ (50)
$F$ $=\frac{|\langle\mathrm{MPS}|\mathrm{CP}\rangle|^{2}}{\langle\mathrm{MPS}|\mathrm{MPS}\rangle\langle\mathrm{CP}|\mathrm{CP}\rangle}.$ (51)
When the CP decomposition exactly reproduces the ground-state wave function, the infidelity becomes zero. For simplicity, all the elements in $|\mathrm{MPS}\rangle$ and $|\mathrm{CP}\rangle$ are assumed to be real in the transverse-field Ising model. The other is the energy difference between the exact ground-state energy $E_{\mathrm{exact}}$ and the energy calculated from the approximated CP decomposed tensor ($E_{\mathrm{CP}}$), defined by
$\Delta E$ $=\left|\frac{E_{\mathrm{CP}}-E_{\mathrm{exact}}}{E_{\mathrm{exact}}}\right|,$ (52)
$E_{\mathrm{CP}}$ $=\frac{\langle\mathrm{CP}|H|\mathrm{CP}\rangle}{\langle\mathrm{CP}|\mathrm{CP}\rangle}.$ (53)
We see that both of these errors decrease polynomially as the rank of the CP decomposition increases.
Figure 4: Infidelity as a function of iterations for $h/J=2$ and $n=16$. We show the results for ranks ranging from $R=8$ to $R=64$. Different colors represent different ranks of the CP decomposition, whereas different symbols represent different initial factors for the CP decomposition.
Figure 5: (a) Infidelity and energy difference as a function of the rank $R$ of the CP decomposition for $h/J=2$ and $n=16$. Different symbols represent different initial factors for the CP decomposition. Open symbols are for the infidelity and filled symbols are for the energy difference. (b) Same as panel (a) but shown on a logarithmic scale. We plot the line proportional to $R^{-2}$ as a reference. Note that, hereafter, all energy values are expressed in the units of $J$.
Figure 6: Infidelity as a function of iterations for $h/J=0.5$ and $n=16$. We show the results for ranks ranging from $R=8$ to $R=64$. Different colors represent different ranks of the CP decomposition, whereas different symbols represent different initial factors for the CP decomposition.
Figure 7: (a) Infidelity and energy difference as a function of the rank $R$ of the CP decomposition for $h/J=0.5$ and $n=16$. Different symbols represent different initial factors for the CP decomposition. Open symbols are for the infidelity and filled symbols are for the energy difference. (b) Same as panel (a) but shown on a logarithmic scale. We plot the line proportional to $R^{-2}$ as a reference.
Figure 8: Infidelity as a function of iterations for $h/J=1$ and $n=16$. We show the results for ranks ranging from $R=8$ to $R=64$. Different colors represent different ranks of the CP decomposition, whereas different symbols represent different initial factors for the CP decomposition.
Figure 9: (a) Infidelity and energy difference as a function of the rank $R$ of the CP decomposition for $h/J=1$ and $n=16$. Different symbols represent different initial factors for the CP decomposition. Open symbols are for the infidelity and filled symbols are for the energy difference. (b) Same as panel (a) but shown on a logarithmic scale. We plot the line proportional to $R^{-2}$ as a reference.
We first examine the model under open boundary conditions for $h/J=2$. The system size is chosen as $n=16$, and the maximum bond dimension is set as $D_{\mathrm{max}}=8$ for the DMRG simulation. For the fixed system size $n$ and the fixed bond dimension $D_{\mathrm{max}}$, we increase the rank of the CP decomposition from $R=8$ to $R=64$ and calculate the errors. Note that the CP decomposition by the ALS method itself does not require the knowledge of the Hamiltonian $H$. The Hamiltonian $H$ is only required to calculate the energy difference between the ground state and the approximated CP decomposed tensor.
As shown in Fig. 4, the infidelity between the ground-state MPS and the approximated CP decomposed tensor almost converges to a small constant value for iterations larger than a few hundred steps. We consider $10$ different sets of random initial factors for the CP decomposition. Regardless of the choice of the initial factors, the infidelity nearly converges to the same value for the same rank of the CP decomposition, suggesting that the CP decomposition is less susceptible to the random initial state. As $R$ increases, the converged value of the infidelity gradually decreases.
We illustrate the $R$ dependence of the errors in Fig. 5. Both infidelity and energy difference decrease monotonically as $R$ increases. The decay appears to be faster than $\sim R^{-1}$. Indeed, for larger $R$, $R$ dependencies of both errors are well approximated by the line proportional to $R^{-2}$.
We then examine the case for $h/J=0.5$. As shown in Fig. 6, similarly to the case for $h/J=2$, the infidelity converges to a small constant value within a few hundred iterations. When the rank of the CP decomposition is fixed, the converged value of the infidelity is typically one digit smaller than the case for $h/J=2$. We also observe that the infidelity is rather sensitive to the initial factor of the CP decomposition.
We also show the $R$ dependence of the errors in Fig. 7. Both infidelity and energy difference decrease rapidly as $R$ increases. When $R\geq 24$, the energy error becomes less than $10^{-4}$ and is almost invisible from the plot. When plotting the errors in the logarithmic scale, we confirm that the decay appears to obey the same $R$ dependence ($\sim R^{-2}$) as is the case for $h/J=2$.
Finally, we examine the case for $h/J=1$, corresponding to the critical point of the model. At the critical point, the correlation length diverges, and the ground-state wave function contains long tails that do not decay to zero at long distances. Such a wave function is generally hard to represent with a small number of parameters. However, we show that the CP decomposition is able to represent the ground-state wave function, just as in the cases for $h/J=0.5$ and $h/J=2$. As shown in Fig. 8, the infidelity again converges to a small constant value within a few hundred iterations. The value is larger than the cases for $h/J=0.5$ and $h/J=2$, but it gradually decreases as $R$ increases. The overall trend is similar to the case for $h/J=2$, and the initial-factor dependence of the CP decomposition is small even at the critical point.
We show the $R$ dependence of the errors at the critical point in Fig. 9. Both infidelity and energy difference decrease monotonically as $R$ increases. For larger $R$, $R$ dependencies of both errors satisfy the power-law decay $\sim R^{-2}$. The exponent of the decay is consistent with what we observed in the cases for $h/J=0.5$ and $h/J=2$. Both infidelity and energy difference behave in the same way for all parameters $h/J$ that we have considered. Therefore, both quantities can be viewed as an equivalent indicator of the errors.
III.2 VMC simulation under open boundary conditions
Figure 10: VMC simulation for $h/J=1$ and $n=16$ under open boundary conditions. The initial state is generated by the CP decomposition of the ground-state MPS with the maximum bond dimension $D_{\mathrm{max}}=8$ under open boundary conditions. With increasing the rank $R$ of the CP decomposition, the energy approaches the true ground-state energy for open boundary conditions.
Figure 11: (a) Energy error as a function of the rank $R$ of the CP decomposition for $h/J=1$ and $n=16$. Filled circles represent the energy error ($|E_{\mathrm{initial,open}}-E_{\mathrm{exact,open}}|/|E_{\mathrm{exact,open}}|$) of the initial state for open boundary conditions used for the VMC method. Up-pointing and down-pointing triangles represent the energy error ($|E_{\mathrm{VMC,open}}-E_{\mathrm{exact,open}}|/|E_{\mathrm{exact,open}}|$, $|E_{\mathrm{VMC,periodic}}-E_{\mathrm{exact,periodic}}|/|E_{\mathrm{exact,periodic}}|$) of the optimized state for open and periodic boundary conditions, respectively. (b) Same as panel (a) but shown on a logarithmic scale. We plot the line proportional to $R^{-2}$ as a reference.
We have observed that the CP decomposition is able to represent the ground-state wave function efficiently with increasing the rank of the CP decomposition. To investigate how close the CP decomposed MPSs (the initial state) is to the ground-state wave function within a given rank of the CP decomposition, we apply the VMC method to further optimize the parameters in the RBM wave function. For each iteration, we typically thermalize the system for $10^{4}$ Monte Carlo steps and evaluate the physical quantities for $10^{4}$ Monte Carlo steps. We optimize the parameters in the RBM wave function for $2000$ iterations and average them for the last $100$ iterations to obtain the final parameters. We then calculate the energy and its statistical error using independent measurements for $32$ bins, each of which contains $10^{5}$ Monte Carlo steps for thermalization and $10^{5}$ Monte Carlo steps for evaluation of the physical quantities.
Figure 10 shows the optimization by the VMC method for $h/J=1$ and $n=16$. For smaller $R$, we observe the further decrease of the energy by optimizing the parameters in the RBM wave function using the VMC method. This result suggests that the initial state obtained by the CP decomposition can be further refined to improve the accuracy of the wave function. With increasing $R$, the energy gain by the VMC method gets smaller. The initial state prepared by the CP decomposition is already close to the ground-state wave function, and the optimization does not change the energy significantly.
We summarize the $R$ dependence of the energy error in Fig. 11. After the further optimization by the VMC method, we can slightly improve the accuracy of the wave function. The energy error after the optimization still decreases monotonically as a function of $R$. Because of the improvement of the accuracy for smaller $R$, the $R$ dependence of the energy error becomes less steeper and appears to be well approximated by the line proportional to $R^{-c}$ with $c\lesssim 2$ being a constant.
III.3 VMC simulation under periodic boundary conditions
Figure 12: VMC simulation for $h/J=1$ and $n=16$ under periodic boundary conditions. Similarly to the case in Fig. 10, the initial state is generated by the CP decomposition of the ground-state MPS with the maximum bond dimension $D_{\mathrm{max}}=8$ under open boundary conditions. With increasing the rank $R$ of the CP decomposition, the energy approaches the true ground-state energy for periodic boundary conditions.
In general, it is hard to prepare the MPS representation of the ground-state wave function under periodic boundary conditions. However, the ground-state wave function should be insensitive to boundary conditions when the system is large enough. Therefore, one may seed the wave function for periodic boundary conditions with the initial wave function for open boundary conditions. We demonstrate that the initial state obtained by the CP decomposition for the open-boundary system can also be efficiently used for the simulation under periodic boundary conditions.
Figure 12 shows the optimization by the VMC method for $h/J=1$ and $n=16$. We use the same initial state as in the case in Fig. 10. We can confirm it from the fact that the energy at the first iteration for periodic boundary conditions is nearly the same as that for open boundary conditions. For all $R$, we observe the further decrease of the energy by optimizing the parameters in the RBM wave function using the VMC method. Although the initial state is not necessarily an accurate representation of the ground-state wave function, it is close enough to the true ground state in a sense that the VMC method can efficiently optimize the parameters. Optimized energies monotonically decrease with increasing $R$ and approach the true ground-state energy for periodic boundary conditions.
We also examine the $R$ dependence of the energy error in Fig. 11. As in the case of open boundary conditions, the energy error after the optimization decreases monotonically as a function of $R$. The improvement under periodic boundary conditions is comparable to or slightly better than that under open boundary conditions. The error is nearly approximated by the line proportional to $R^{-c}$ with $c\lesssim 2$ being a constant.
III.4 Size dependence of the required rank of the CP decomposition
Figure 13: Size dependence of the ground-state energy per site for $h/J=1$ under periodic boundary conditions. The rank $R$ of the CP decomposition, which gives $2nR$ variational parameters ($b_{l,k}$ and $W_{l,k}$) in the RBM wave function for the VMC method, is chosen to be $R=n/2$ for all systems of $n$ sites. Circles and triangles represent the exact and VMC energy per site, respectively. As references, we also plot the asymptotic behavior of the energy per site by a dashed line and that multiplied by $0.98$, corresponding to a $2\%$ error, by a dotted line. The energy error obtained by the VMC method is nearly independent of the system size and is less than $2\%$ for all systems.
Figure 14: Size dependence of the ground-state energy per site for $h/J=0.5$ under periodic boundary conditions. The rank $R$ of the CP decomposition is chosen to be $R=n/2$ for all systems of $n$ sites. Circles and triangles represent the exact and VMC energy per site, respectively. As references, we also plot a dashed line interpolating the exact data points and the line multiplied by $0.999$, corresponding to a $0.1\%$ error, by a dotted line.
Figure 15: Size dependence of the ground-state energy per site for $h/J=2$ under periodic boundary conditions. The rank $R$ of the CP decomposition is chosen to be $R=n/2$ for all systems of $n$ sites. Circles and triangles represent the exact and VMC energy per site, respectively. As references, we also plot a dashed line interpolating the exact data points and the line multiplied by $0.96$, corresponding to a $4\%$ error, by a dotted line.
Finally, we numerically investigate how large the rank $R$ of the CP decomposition should be for arbitrary system sizes. We mainly focus on a rather difficult case at the critical point ($h/J=1$) of the transverse-field Ising model under periodic boundary conditions. The ground state is known to exhibit a logarithmic correction in the entanglement entropy, which naturally requires slightly larger bond dimensions in the MPS representation compared to bond dimensions required for conventional gapped ground states. As we see below, for all systems of $n$ sites, the condition $R=\mathcal{O}(n)$ is sufficient to obtain the ground-state wave function having a small energy error that is independent of the system size.
We first prepare the MPS representation of the ground-state wave function with the maximum bond dimension $D_{\mathrm{max}}=n/2$ for all systems of $n$ sites using the DMRG method. The energy error from the true ground state is found to be less than $10^{-9}$ in the units of $J$ for any system size that we consider. We then CP decompose the MPS representation using the ALS method with $100$ iterations from the random initial CP factors. We optimize the parameters in the RBM wave function generated by the CP decomposed tensor using the VMC method. We keep the rank $R$ of the CP decomposition to $R=D_{\mathrm{max}}(=n/2)$ for all systems of $n$ sites. The number of variational parameters ($b_{l,k}$ and $W_{l,k}$) is $2nR=n^{2}$, which is much smaller than the dimension of the Hilbert space. We use $10^{6}$ ($4\times 10^{6}$) Monte Carlo steps for evaluating physical quantities during the optimization (during the statistical processing).
Figure 13 shows the ground-state energy per site for $h/J=1$ under periodic boundary conditions. The exact energy per site ($e_{\mathrm{exact}}$) follows the asymptotic behavior [104, 100, 105, 106, 107], given by
$e_{\mathrm{exact}}(n)=-\frac{4}{\pi}-\frac{\pi}{6n^{2}}+\mathcal{O}(n^{-3}),$ (54)
and the energy per site obtained by the VMC method nearly shows the same behavior. The difference between the exact and VMC energy per site is nearly independent of the system size and is always less than $2\%$. Consequently, we numerically confirm that the condition $R=n/2$ is sufficient to reproduce the ground-state wave function having a small energy error that does not depend on the system size.
Away from the critical point ($h/J=1$), the energy error obtained by the VMC method is also independent of the system size $n$ when the condition $R=\mathcal{O}(n)$ is satisfied. We show the ground-state energy per site with $R=n/2$ for $h/J=0.5$ and $h/J=2$ under periodic boundary conditions in Figs. 14 and 15. The VMC method combined with the initial wave function generated by the CP decomposed MPSs performs effectively regardless of whether the system is at the critical point.
In general, at the critical point, the size dependence of the entanglement entropy contains a logarithmic correction, which requires more parameters to accurately represent the corresponding wave function during numerical simulations. Therefore, one may expect the largest energy error at the critical point ($h/J=1$). However, our numerical results suggest that the relative energy error is slightly larger for a parameter region which is away from the critical point. We do not have a clear explanation for the origin of this behavior; however, the RBM wave function does not prevent the efficient representation of quantum states exhibiting entanglement growth that exceeds area-law behavior. Because of the flexible structure, the RBM wave function does not necessarily suffer from the same limitations as the MPS wave function at corresponding points in the parameter space. These results are actually encouraging, because our approach may enable successful simulations in parameter regions that are typically difficult to handle with the MPS wave function.
IV Summary and outlook
We proposed a method for approximately converting the MPS representation of the ground state in quantum many-body systems into the RBM wave function consisting of a single multinomial hidden unit. This procedure helped us prepare suitable initial states for the VMC simulation using NNQSs. The computational complexity of the method for converting MPSs into RBM wave functions using the CP decomposition scaled polynomially with the number of variational parameters and sites in the system. The error of the initial wave function typically decreases as $\sim R^{-2}$ with increasing the rank $R$ of the CP decomposition. We examined the applicability of the method by taking the transverse-field Ising model as an example and found that the RBM wave function with a single hidden unit well approximated the ground state of the quantum many-body system. Even when the energy of the initial state was slightly away from that of the true ground state, the VMC method was able to efficiently optimize the parameters in the RBM wave function and provided a better energy estimate after the optimization, as we confirmed in the case of periodic boundary conditions. This observation suggests that there remains a high likelihood of obtaining an accurate ground state through the VMC method even if the DMRG simulation is inaccurate, causing the corresponding initial state to be slightly away from the ideal state. We also numerically investigated the required rank $R$ of the CP decomposition for arbitrary system sizes. We empirically found that the condition $R=\mathcal{O}(n)$ is sufficient for any $n$ sites to obtain the ground-state wave function having a small energy error (less than $2\%$ for $R=n/2$ at the critical point) that does not depend on the system size.
Our approach can be extended to more general NNQSs, such as those combining the multinomial RBM $\Psi_{\mathrm{mRBM}}(\{v_{i}\})$] and conventional binomial RBM $\Psi_{\mathrm{bRBM}}(\{v_{i}\})$] wave functions:
$\Psi(\{v_{i}\})$ $=\Psi_{\mathrm{mRBM}}(\{v_{i}\})\Psi_{\mathrm{bRBM}}(\{v_{i}\}),$ (55)
$\Psi_{\mathrm{mRBM}}(\{v_{i}\})$ $=\sum_{j=1}^{R}\exp\left[\sum_{i=1}^{n}\left(b_{i,j}+W_{i,j}v_{i}\right)\right],$ (56)
$\Psi_{\mathrm{bRBM}}(\{v_{i}\})$ $=\exp\left(\sum_{i=1}^{n}a^{\prime}_{i}v_{i}\right)\prod_{j=1}^{n_{h}}2\cosh\left(b^{\prime}_{j}+\sum_{i=1}^{n}W^{\prime}_{i,j}v_{i}\right),$ (57)
where $b_{i,j}$, $W_{i,j}$, $a^{\prime}_{i}$, $b^{\prime}_{j}$, and $W^{\prime}_{i,j}$ are variational parameters. The binomial RBM wave function further improves the approximation of the ground state in quantum many-body systems. The conventional binomial RBM wave function would efficiently represent quantum states that exhibit volume-law entanglement entropy [43, 47]. While the multinomial RBM wave function may struggle to capture sufficient entanglement entropy needed for accurately modeling quantum states in some unfortunate cases, combining the binomial RBM wave function can offer a more effective alternative. The integrated RBMs are unlikely to significantly hinder the variational wave function’s convergence toward the ground state when the initial state is already close to the target state.
In strongly correlated electron systems, the low-energy eigenstates often exhibit complex nodal structures [108]. In most of the VMC simulations, we imitate such nodal structures by using one-body wave functions constructed from the Slater determinant or the Pfaffian, derived from ground states of noninteracting fermion systems or mean-field solutions of Fermi-Hubbard systems [109, 110]. The CP decomposition has recently been employed to enhance fermionic trial wave functions within the VMC method [111]. Our method would provide a complementary way to prepare wave functions with complex nodal structures, which will be more accurate starting points than the one-body wave functions consisting of the Slater determinant or the Pfaffian, once we have accurate tensor networks representing the low-energy eigenstates in strongly correlated electron systems.
We have thoroughly examined the performance of the proposed method in the simplest quantum one-dimensional systems to deepen the understanding of how one can efficiently transform tensor network states into shallow RBM wave functions. It is an interesting future direction to extend our method to higher-dimensional systems. In higher dimensions, optimized tensor network states could be inaccurate, causing the corresponding initial RBM wave functions to be far from the ground state. Nevertheless, this initial energy is expected to be much closer to the ground-state energy than those obtained from random initializations and those prepared by the conventional mean-field approximation. Then, the optimization from the initial RBM wave function prepared by our method is more likely to be successful. Investigating the performance of our method in general higher-dimensional systems requires much more careful analyses and will be left for future work.
In the present study, we mainly focused on wave functions that explicitly break the lattice translational symmetry. In translational invariant systems, we may use the translational invariant RBM wave function that has a much smaller number of variational parameters. Such wave functions should be generated using the symmetric CP decomposition [112] of translational invariant tensor networks, including the infinite MPS (iMPS) [3] in one-dimensional systems and the infinite PEPS (iPEPS) [11, 12, 9, 13, 10, 14] in two-dimensional systems. One may apply the ALS method to iMPS or iPEPS to obtain translationally invariant factors for the CP decomposed tensor networks. It will be interesting to see to what extent the CP decomposition of iMPS or iPEPS helps us prepare suitable initial states for the translational invariant RBM wave function. However, this is beyond the scope of this paper and is left for future work.
Note that, even for systems that break the translational symmetry, tensor network states in higher dimensions possess structural and computational characteristics that may differ significantly from the present shallow RBM wave function and the 1D tensor network state. Therefore, generalizing our method to higher dimensions is not straightforward, and one may develop a more specific method for converting PEPSs into RBM wave functions. Moreover, the optimization of PEPS itself remains a challenging and active research field. Addressing these challenges is an interesting future direction.
Finally, we would like to point out that the CP decomposition of MPSs and tensor trains [1, 2, 3, 4, 5, 6, 7, 8, 9, 10] is a general technique that can be applied to other research fields, such as information science and engineering [113, 114]. One may consider the Monte Carlo approach to obtain the CP decomposition of MPSs or tensor trains using the RBM-type neural network. Indeed, extended Boltzmann machine representations of the CP decomposition have been proposed for non-negative tensors, formulated as a convex optimization problem based on information geometry recently [115, 116]. Such approach is also an interesting future direction.
Acknowledgements.
This work was financially supported by MEXT KAKENHI, Grant-in-Aid for Transformative Research Area (Grants No. JP22H05111 and No. JP22H05114). R.K. was supported by JSPS KAKENHI (Grants No. JP24H00973 and No. JP25K07157). S.G. was supported by JSPS KAKENHI (Grant No. JP23K13042). The numerical computations were performed on computers at the Yukawa Institute Computer Facility, Kyoto University and on computers at the Supercomputer Center, the Institute for Solid State Physics, the University of Tokyo.
Data availability
The data that support the findings of this article are openly available [117].